# Latent BKI: Open-Dictionary Continuous Mapping in Visual-Language Latent Spaces with Quantifiable Uncertainty

### v1

@@ III-C Inference
After updating the map, we can compute an expectation and variance for features observed within each voxel. First, the distribution can be marginalized to obtain a posterior predictive solution that defines the probability of observing a feature $y_{*}$ at voxel centroid $x_{*}$. The posterior predictive distribution for voxel * is:
$p(y_{*}|x_{*})=t_{\lambda_{*}}\left(\mu_{*},\frac{\lambda_{*}+1}{\lambda_{*}^{2}}\Psi_{*}\right),$ (16)
where $t_{\lambda_{*}}$ is the multivariate Student-$t$ distribution. The multivariate Student-$t$ distribution has an expectation of: $\mathbb{E}(y_{*})=\mu_{*}$, when $\lambda_{*}>1$, and a variance of: $\text{Var}(y_{*})=\frac{\lambda_{*}}{\lambda_{*}-2}\Psi_{*}$, for $\lambda_{*}>2$. However, both the expectation and variance are within the latent space of the neural network and must be decoded to obtain a meaningful interpretation.
@@ III-D Uncertainty Quantification
While the map update step described above can propagate uncertainty, the variance is limited to the latent space of the neural network encoder. Therefore, we propose two methods to quantify uncertainty.
@@
Based on a common approach of information quantification in optimal experimental design [40], we can quantify uncertainty through p-optimality [41]. Although p-optimality leads to many solutions [42], in this work, we compute the uncertainty $U_{*}$ via E-optimality criterion as:
$U_{*}=\max\left(\text{Var}(y_{*})\right),$ (18)
and the uncertainty via D-optimality criterion as:
$U_{*}=\exp\left(\frac{\sum_{i=1}^{C}\log\left(\text{Var}(y_{*})_{i}\right)}{C}\right),$ (19)
where a large value corresponds to high uncertainty. This is because, experimentally, we find that E-optimality is highly correlated with the sampling-based uncertainty and is quick to compute. See Section IV-C for detailed experiments.
@@ IV Results and Discussion
Indoor Segmentation 66.20 20.95 N/A
ConvBKI (Single) 68.87 24.28 Fixed
LatentBKI 68.16 23.14 Open
Outdoor Segmentation 90.46 52.05 N/A
ConvBKI (Single) 91.01 56.38 Fixed
LatentBKI 91.29 57.33 Open
@@ IV-A Comparison against categorical space mapping method
First, we compare LatentBKI against the closed-dictionary semantic mapping method defined in [29], which leverages BKI with a categorical likelihood to update the map. Specifically, we compare against ConvBKI [25] using a single un-trained spherical kernel for direct comparison. Our goal of this study is to verify that LatentBKI can generalize BKI into the latent space of neural networks without any significant changes in performance.
We evaluate LatentBKI and ConvBKI (single) on an indoor and outdoor dataset. For the indoor dataset, we use Matterport3D (MP3D) [43] with an image semantic segmentation network, Language-driven Semantic Segmentation (LSeg) [7]. For the outdoor dataset, we choose the Semantic KITTI dataset [44] and use the Sparse Point-Voxel Convolution Neural Network (SPVCNN) [45, 46] for point cloud semantic prediction. Since the latent space of LSeg has a large dimension of 512, we use Principal Component Analysis (PCA) to down-sample features for the map update. The PCA algorithm is trained on held-out MP3D sequences we do not compare with ConvBKI on, and we experimentally find that reducing the size of the latent space of LSeg to 64 results in only a $0.09\%$ loss in accuracy and $0.42\%$ decrease in mIoU.
We apply the same configuration to both mapping algorithms to ensure comparable results. Each algorithm uses a voxel resolution of 0.1 meters, a sparse kernel with a kernel length of 0.5 meters, and a filter size of 3, determining how many neighboring voxels should be updated for a single-point observation. We apply a cubic filter in this case, so $273^{3}=$ 27 neighboring voxels will be considered. Since both methods compare spatial smoothing, we provide 80% of the points as input and evaluate semantic predictions over the mean intersection over union (mIoU) and accuracy metrics on the remaining 20% of the points.
As shown in Table I, LatentBKI performs similarly to ConvBKI over indoor and outdoor datasets. These results verify that BKI with a Gaussian likelihood is suitable for the latent space of multiple neural networks and does not cause any significant changes in performance. While LatentBKI results in a marginal improvement in quantitative performance on the outdoor data, the slight decrease in indoor data is due to the dimensionality reduction applied by PCA on the input to LatentBKI. Additionally, although the studies are performed on closed-dictionary segmentation, LatentBKI enables open-dictionary continuous mapping as the per-voxel latent features can be decoded into more meaningful outputs, which we demonstrate in Section IV-D. Next, we compare LatentBKI with a popular latent mapping algorithm, VLMap.
@@ IV-B Latent Mapping Comparison
We compare LatentBKI with a similar open-dictionary mapping method VLMap [36], which also updates voxels through a weighted averaging approach. Since VLMap does not utilize spatial information, we compare LatentBKI to VLMap with a filter size of 1, disregarding points in nearby voxels when calculating the posterior state of a voxel. Both methods use the open-dictionary encoder LSeg, with features downsampled by PCA to a latent size of $64$.
We compare each method on the MP3D dataset [43] following the same experimental setup as VLMap, including a resolution of $0.05$ m to account for the fine-resolution indoor environment. To ensure that the input and evaluation points are consistent, we implemented the pre-processing heuristics of VLMap within our method. These heuristics include discarding pixels with extreme depths, discarding points outside of the scene range, and downsampling input points. Specifically, the same set of $1\%$ of remaining pixels are provided to each method, and points with depth less than $0.1$ meters or greater than $6$ meters are excluded. By following the same downsampling heuristics as VLMap in our evaluation, we isolate the effect of the sparse kernel function used by our method to weight input points compared to the depth-wise weighting scheme employed by VLMap.
The results of our experiments on two different MP3D sequences, as shown in Table II, indicate that LatentBKI outperforms VLMap in both accuracy and mean IoU, which we attribute to the sparse kernel. At its core, the VLMap algorithm is similar to our method of calculating the expectation with a different weighting scheme, and without spatial smoothing or quantifiable uncertainty. This performance improvement may be attributed to the formalized mathematical updates in LatentBKI, which provide a more robust integration of observations and better handling of uncertainty. Additionally, the formal probabilistic approach we leverage in LatentBKI allows for quantifiable uncertainty, which we study next.
Table II: Comparison against VLMap, which performs latent mapping without spatial smoothing or uncertainty quantification. Results are compared on scenes gTV8FGcVJC9 (1) and 5LpN3gDmAk7 (2) of MP3D.
@@
1 Segmentation 61.08 19.57 N/A
VLMap 60.94 20.32 No
LatentBKI 63.30 21.56 Yes
2 Segmentation 63.27 13.12 N/A
VLMap 64.12 12.44 No
LatentBKI 66.10 14.92 Yes
@@ IV-C Ablation Studies
Spatial Smoothing: In real-world applications, data is often sparse due to sensors such as LiDAR or sparse stereo matching algorithms. BKI provides a probabilistic technique to create more complete maps from sparse spatial data by leveraging the spatial smoothing effect of kernels. In this experiment, we compare different kernel sizes and their ability to complete the map from sparse data.
All kernels are compared on the same scene of MP3D, where data is downsampled temporally to incorporate only one in 3 frames and at the image level to use a randomly sampled set of pixels from each image. Fig. 3 demonstrates plots of the segmentation performance of different kernel sizes and varying sparsity levels. At extreme sparsity levels, spatial smoothing of points benefits the map completeness. As the input becomes more dense, in this case, due to high-resolution images with accurate depth, the effect of the kernel is diminished.
@@
Uncertainty Quantification: To evaluate the ability of LatentBKI to quantify uncertainty meaningfully, we quantitatively and qualitatively compare uncertainty quantification on the MP3D dataset using LSeg as the encoder. We construct a map using LatentBKI, then compare uncertainty quantified using D-optimality, E-optimality, and sampling as described in Section III-D. Whereas the sampling-based method is commonly used, it is computationally expensive compared to the E-optimality and D-optimality-based techniques.
@@ V Conclusion
While LatentBKI demonstrated success in open-dictionary environments through a Gaussian likelihood, there are several avenues for future work. First, LatentBKI does not consider the unique geometry of objects and can therefore be combined with architectures such as ConvBKI, which learns per-category kernels, or the high-quality 3D Gaussian Splatting [48] novel view synthesis method which represents the environment using 3D ellipsoids, similar to the kernel structure we leverage for continuous mapping. Additionally, inspired by the recent work on open-dictionary radiance fields [49], we believe that the segment anything model [4, 5] may be useful when identifying the boundaries of objects.

### v2

@@ III-C Inference
After updating the map, we can compute an expectation and variance for features observed within each voxel. First, the distribution can be marginalized to obtain a posterior predictive solution that defines the probability of observing a feature $\bm{y}_{*}$ at voxel centroid $\bm{x}_{*}$. The posterior predictive distribution for voxel * is:
$p(\bm{y}_{*}|\bm{x}_{*})=t_{\lambda_{*}}\left(\bm{\mu}_{*},\frac{\lambda_{*}+1}{\lambda_{*}^{2}}\Psi_{*}\right),$ (12)
where $t_{\lambda_{*}}$ is the multivariate Student-$t$ distribution. The multivariate Student-$t$ distribution has an expectation of: $\mathbb{E}(y_{*})=\bm{\mu}_{*}$, when $\lambda_{*}>1$, and a covariance of: $\text{Cov}(\bm{y}_{*})=\frac{\lambda_{*}}{\lambda_{*}-2}\left(\frac{\lambda_{*}+1}{\lambda_{*}^{2}}\Psi_{*}\right)$, for $\lambda_{*}>2$. However, both the expectation and covariance are within the latent space of the neural network and must be decoded to obtain a meaningful interpretation.
@@ III-D Uncertainty Quantification
While the map update step described above can propagate uncertainty, the covariance is limited to the latent space of the neural network encoder. Therefore, we propose two methods to quantify uncertainty.
@@
Based on a common approach of information quantification in optimal experimental design [25], we propose to consolidate the covariance matrices as a single scalar value $U_{*}$ through p-optimality [34]. Although p-optimality leads to many solutions [35], in this work we calculate and compare E-optimality, which selects the maximum eigen-value of the covariance matrix, and D-optimality, which calculates the volume of the covariance hyper-ellipsoid. For eigen-values $\bm{\lambda}$ of the covariance matrix, which we note are the diagonals of an uncorrelated covariance matrix, E-optimality criterion is computed as: $U_{*}=\max\left(\bm{\lambda}\right)$, and D-optimality criterion is computed as:
$U_{*}=\exp\left(\frac{\sum_{i=1}^{C}\log\left(\bm{\lambda}_{i}\right)}{C}\right).$ (14)
Experimentally, we find that E-optimality is highly correlated with the sampling-based uncertainty and is quick to compute. See Section IV-C for detailed experiments
III-E Feature Compression
Due to the large latent dimension of VLM’s, we propose to make two approximations to reduce computation and memory complexity. First, we approximate the covariance matrix with only the diagonal elements, significantly reducing complexity at the cost of cross-correlation terms. Diagonal covariances are common in the feature space and are used in variational auto-encoders (VAEs). Second, we use PCA to reduce the latent dimension of encoded features from a latent dimension of 512 to 64 from VLM’s before fusing into our map. PCA is an unsupervised learning algorithm that maximizes information preserved during compression and uses an affine transformation to upscale the compressed features back to the original dimension. Since the transformation is affine, we can compute the full dimensional expectation by passing the compressed expectation through the PCA upsampling. For the uncertainty, we compute E and D optimality in the reduced latent space and obtain full dimensional samples for the sampling technique through PCA upsampling. Overall, we found that PCA is generalizable to new scenes with minimal performance loss from compression.
@@ IV Results and Discussion
Quantitative results are obtained on popular outdoor and indoor datasets. For the outdoor results, we compare on the validation set of the Semantic KITTI dataset [36] using the Sparse Point-Voxel Convolution Neural Network (SPVCNN) [37] for point cloud semantic prediction. We choose the validation set because it has publicly available ground truth, and $4,070$ frames. For the indoor comparison, we evaluate methods on all eight scenes of the Matterport3D (MP3D) [38] dataset with an open-dictionary image semantic segmentation network, Language-driven Semantic Segmentation (LSeg) [8]. Since the latent space of LSeg has a large dimension of 512, we use PCA to down-sample features for the map update to a size of 64.
@@
Indoor Segmentation 59.14 14.64 N/A
ConvBKI (Single) 61.49 16.69 Fixed
LatentBKI 60.44 16.15 Open
Outdoor Segmentation 89.60 58.54 N/A
ConvBKI (Single) 90.02 61.26 Fixed
LatentBKI 90.02 61.54 Open
@@ IV-A Comparison against categorical space mapping method
First, we compare LatentBKI against the closed-dictionary semantic mapping method defined in [16], which leverages BKI with a categorical likelihood to update the map. Specifically, we compare against ConvBKI [4] using a single un-trained spherical kernel for direct comparison. Our goal of this study is to verify that LatentBKI can generalize BKI into the latent space of neural networks without any significant changes in performance. Note that in this experiment, similar quantitative results indicate successful application of BKI to the latent space without any loss of functionality.
We apply the same configuration to both mapping algorithms to ensure comparable results. Each algorithm uses a voxel resolution of 0.1 meters, a sparse kernel with a kernel length of 0.5 meters, and a filter size of 3, determining how many neighboring voxels should be updated for a single-point observation along a single axis. Since both methods compare spatial smoothing, we provide 80% of the points as input and evaluate semantic predictions over the mean intersection over union (mIoU) and accuracy metrics on the remaining 20% of the points.
As shown in Table I, LatentBKI performs similarly to ConvBKI over indoor and outdoor datasets without any decrease in performance. These results verify that our approach generalizes BKI to the latent space of networks, enabling open-dictionary probabilistic mapping, successfully. While LatentBKI results in a marginal improvement in quantitative performance on the outdoor data, the slight decrease in indoor data is due to the dimensionality reduction applied by PCA on the input to LatentBKI. Next, we compare LatentBKI with a popular latent mapping algorithm, VLMap.
@@ IV-B Latent Mapping Comparison
We compare LatentBKI with a similar open-dictionary mapping method VLMap [10], which also updates voxels through a weighted averaging approach. We choose to compare specifically with VLMap since weighted averaging is a popular technique for open-dictionary mapping [13, 11, 10], and the volumetric representation of VLMap allows for a direct and conclusive comparison. Since our approach generalizes VLMap to include a spatial kernel and quantifiable uncertainty, we compare the results with ($k=3$) and without ($k=1$) spatial smoothing, where $k$ is the number of neighboring voxels along each dimension an observation can influence. To demonstrate the benefits of VLMap, we also implement a heuristic baseline that stores the feature of the most recent coinciding observation within each voxel.
We compare each method on the MP3D dataset [38] following the same experimental setup as VLMap, including a resolution of $0.05$ m to account for the fine-resolution indoor environment. Additionally, following the setup of VLMap we discard pixels with extreme depths $<0.1$ m or $>6$ m, discard points outside of the scene range, and downsample input points to the same set of $1\%$ of input pixels. By following the same downsampling heuristics as VLMap in our evaluation, we isolate the effect of the sparse kernel function and spatial smoothing used by our method to weight input points compared to the depth-wise weighting scheme employed by VLMap [13, 39].
The results of our experiments in Table II indicate that both LatentBKI and VLMap outperform the heuristic baseline, demonstrating the benefit of the weighted average approach. LatentBKI outperforms VLMap in both accuracy and mean IoU, which we attribute to the sparse kernel and spatial smoothing. We note that our method is a probabilistic generalization of VLMap with a spatial kernel and quantifiable uncertainty, however these benefits also increase computational complexity linearly with the size of the spatial kernel. Although the computational complexity of our method is greater, our implementation is more efficient than VLMap due to vectorization and the use of the GPU, requiring $122.05$ ms to update $100,000$ points compared to $4,325.74$ ms for VLMap to update the same number of points.
Table II: Latent mapping comparison on MP3D.
@@
Segmentation [8] 53.24 12.59 N/A
Heuristic 51.63 11.60 No
VLMap [10] 53.84 12.53 No
LatentBKI ($k=1$) 55.57 14.01 Yes
LatentBKI ($k=3$) 55.86 14.18 Yes
@@ IV-C Ablation Studies
Spatial Smoothing: In real-world applications, data is often sparse due to sensors such as LiDAR or sparse stereo matching algorithms. BKI provides a probabilistic technique to create more complete maps from sparse spatial data by leveraging the spatial smoothing effect of kernels. In this experiment, we compare different kernel sizes ($k$) and their ability to complete the map from sparse data.
All kernels are compared on the same scene of MP3D, where data is downsampled temporally to incorporate only one in 3 frames and at the image level to use a randomly sampled set of pixels from each image. Fig. 3 demonstrates plots of the segmentation performance of different kernel sizes and varying sparsity levels. At extreme sparsity levels, spatial smoothing of points benefits the map completeness. As the input becomes more dense, the effect of the kernel is diminished.
@@
Uncertainty Quantification: To evaluate the ability of LatentBKI to quantify uncertainty meaningfully, we quantitatively and qualitatively compare uncertainty quantification on the MP3D dataset using LSeg as the encoder. We construct a map using LatentBKI, then compare uncertainty quantified using D-optimality, E-optimality, and sampling as described in Section III-D. Whereas the sampling-based method is commonly used, it is computationally expensive compared to the E-optimality and D-optimality-based techniques with a run-time of $5,661$ ms for $10,000$ query voxels compared to $2.3$ ms for the optimality approaches to compute.
@@ V Conclusion
While LatentBKI demonstrated success in open-dictionary environments through a Gaussian likelihood, there are several avenues for future work. First, LatentBKI does not consider the unique geometry of objects and can therefore be combined with architectures such as ConvBKI, which learns per-category kernels, or the high-quality 3D Gaussian Splatting [41] novel view synthesis method which represents the environment using 3D ellipsoids, similar to the kernel structure we leverage for continuous mapping. Additionally, inspired by the recent work on open-dictionary radiance fields [42], we believe that the segment anything model [5] may be useful when identifying the boundaries of objects. Last, the uncertainty produced by LatentBKI may be used for planning tasks such as active perception by quantifying expected information gain [18, 19, 25].